OpenStudy (anonymous):

A right triangle has a hypotenuse of 25 feet and a leg of 7 feet. Which of these is the length of the other leg? 12 feet 15 feet 21 feet 24 feet

OpenStudy (johnweldon1993):

 Lets solve for that other side. Since this is a right triangle we can use the Pythagorean Theorem. \[\large a^2 + b^2 = c^2\] where c = hypotenuse of the triangle and a, b are the other 2 legs... We have the hypotenuse...and let say we have 'a'....so we need to rearrange his equation to solve for 'b' \[\large b^2 = c^2 - a^2\] Well we want 'b' not b^2...so take the square root of both sides \[\large b = \sqrt{c^2 - a^2}\] Now plug in your 25 for 'c' and 7 for 'a' and you can solve for 'b'

OpenStudy (anonymous):

D